OpenStudy (anonymous):

twice the sum of the square of a number and 1, increased by the square of the quantity 1 more than the square of the number gives 3. Find the numbers. Answers are 0,+- 2i

OpenStudy (pawanyadav):

2(x^2+1)+(1+x^2)^2=3

OpenStudy (pawanyadav):

Let 1+x^2=t So. t^2+2t=3 t^2+2t-3=0 You will get t=1,-3 Put 1+x^2=1 x^2=1-1=0 So x=0 Now put 1+x^2=-3 x^2=-2 x=+-2i So we get three values of x 0,+2i,-2i

OpenStudy (pawanyadav):

Any appreciation for this help

OpenStudy (whpalmer4):

\[2(x^2+1)+(x^2+1)^2=3\] \[2x^2+2+x^4+2x^2+1=3\]\[x^4+4x^2+3=3\]\[x^4+4x^2=0\]\[x^2(x^2+4)=0\]\[x^2=0\]that gives us our first answer of \(x = 0\)\[x^2+4=0\]\[x^2=-4\]\[\sqrt{x^2} = \sqrt{-4}\]\[\sqrt{x^2} = \sqrt{4*-1}\]\[\sqrt{x^2} = 2\sqrt{-1}\]\[x=\pm2i\]for our other two answers

OpenStudy (whpalmer4):

@pawanyadav you made a small algebra error: Now put 1+x^2=-3 x^2=-2 x=+-2i that should be \[1+x^2=-3\]\[1-1+x^2 = -3-1\]\[x^2=-4\]\[x = \pm 2i\]

OpenStudy (pawanyadav):

Sorry wrongly typed